OpenStudy (anonymous):

The mole to mole ratio of hydrogen peroxide to oxygen is 2:1. Use this and 5.56 x 10 ^23 molecules of hydrogen peroxide present to determine the liters of oxygen produced at room temp.

OpenStudy (anonymous):

*@RTP= 25.4

OpenStudy (anonymous):

ok so just use: n=N/L where : n - moles of atoms/molecules N - number of atoms/molecules L - 6,022 * 10^23 mol-1 and then: n(H2O2)/n(O2) = 2/1 n(H2O2) = 2*n(O2) and then: n=V/Vm V= n*Vm where V - volume n - number of moles Vm - as you wrote 25,4 dm3mol-1

OpenStudy (anonymous):

Can you help me through it?

OpenStudy (anonymous):

just put numbers in those equations...

OpenStudy (anonymous):

For the first one, L, wouldnt it be 6.02 x 10^23?

OpenStudy (anonymous):

yes